ECH 4604: Final Exam Fall 2004

1. Choose a design problem from the unit operations laboratory, the measurements laboratory, separations, thermodynamics, heat transfer or fluid mechanics courses.

2. Make sure that this problem cannot be solved by the CHEMCAD modules available for the design course.

3. Develop a module along the lines available in the Design I class that solves this problem using CHEMCAD.

4. This problem is due on Thursday, December 9, 2004 at 10:00 a.m. (this is the scheduled date and time for the final for this course).

Please note the following:

• You need to get approval from the course instructor by the last day of class for the problem that you intend to simulate in CHEMCAD .

• This exam is worth 3 points. If you do not turn in your module by the specified date and time, you will get a zero for this exam.

• This is not a team assignment; each student is expected to work on this problem individually and the usual honor code rules for exams apply to this assignment.